# Measurement of the branching fraction of $\eta\rightarrow\mu^{+}\mu^{-}$ and search for $\eta\rightarrow e^{+}e^{-}$

### v1

@@ VII Summary
With a sample of (10087 $\pm$ 44) $\times$ 106  BESIII:2021cxx $J/\psi$ events collected with the BESIII detector, we perform a study of $\eta\to\ell^{+}\ell^{-}$ tagging with $\eta^{\prime}\to\pi^{+}\pi^{-}\eta$ in $J/\psi$ radiative decays. The decay $\eta\rightarrow\mu^{+}\mu^{-}$ is observed with significance of $9.8\sigma$, and its decay branching fraction is measured to be $(5.8\pm 1.0_{\rm stat}\pm 0.2_{\rm syst})\times 10^{-6}$. This is consistent with theoretical predictions Petri:2010ea and the $(5.8\pm 0.8)\times 10^{-6}$ PDG value PhysRevD.110.030001. No signal from the $\eta\rightarrow e^{+}e^{-}$ is observed, and the upper limit on its decay branching fraction at 90% CL is set at $2.2\times 10^{-7}$. The result improves the upper limit of $7\times 10^{-7}$ $(CL=90\%)$ from the SND experiment SND:2018egu, providing a stricter experimental constraint on new physics.
@@
In the future, new experimental facilities, such as the super $\tau$-charm facility (STCF) Achasov_2023, the JLab Eta Factory (JEF) Somov:2024jiy, REDTOP Gatto:2019dhj and the $\eta$ factory of HIAF An:2025lws experiments, will provide larger $\eta$ samples. They are expected to conduct precise studies of the $\eta\rightarrow e^{+}e^{-}$ decay, and thereby deeply explore the internal structure of the $\eta$ meson, and search for new physics phenomena beyond the SM.
VIII Acknowledgement
@@
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017,12475089 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; ERC under Contract No. 758462; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; STFC (United Kingdom); Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374

### v2

@@ VII Summary
With a sample of (10087 $\pm$ 44) $\times$ 106  Ablikim and others (2022) $J/\psi$ events collected with the BESIII detector, we perform a study of $\eta\to\ell^{+}\ell^{-}$ tagging with $\eta^{\prime}\to\pi^{+}\pi^{-}\eta$ in $J/\psi$ radiative decays. The decay $\eta\rightarrow\mu^{+}\mu^{-}$ is observed with a significance of $9.8\sigma$, and its branching fraction is measured to be $(5.8\pm 1.0_{\rm stat}\pm 0.2_{\rm syst})\times 10^{-6}$. This is consistent with theoretical predictions Petri (2010) and the $(5.8\pm 0.8)\times 10^{-6}$ PDG value Navas and others (2024). No signal is observed for $\eta\rightarrow e^{+}e^{-}$ , and the upper limit on its branching fraction at 90% CL is set to $2.2\times 10^{-7}$. This represents an improvement over the previous upper limit of $7\times 10^{-7}$ $(CL=90\%)$ from the SND experiment Achasov and others (2018), providing a stricter experimental constraint on new physics. Recently, theoretical predictions based on a dispersive representation for the $\eta$ transition form factors give $\mathcal{B}(\eta\to\mu^{+}\mu^{-})=(4.54\pm 0.04\pm 0.02)\times 10^{-6}$ and $\mathcal{B}(\eta\to e^{+}e^{-})=(5.37\pm 0.04\pm 0.02)\times 10^{-9}$ Messerli et al. (2025). Our measured branching fraction for $\eta\to\mu^{+}\mu^{-}$ is in good agreement with this prediction, with the difference being well within $1.2\,\sigma$.
@@
In the future, new experimental facilities, such as the super $\tau$-charm facility (STCF) Achasov and others (2023), the JLab Eta Factory (JEF) Somov (2024), REDTOP Gatto ,and the $\eta$ factory of HIAF An and others experiments, will provide larger $\eta$ samples, enabling improved measurements of $\eta\rightarrow\ell^{+}\ell^{-}$ decays, deeper exploration of the internal structure of the $\eta$ meson, and searches for new physics beyond the SM.
Acknowledgements
@@
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2025YFA1613900, 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017,12475089 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; ERC under Contract No. 758462; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; STFC (United Kingdom); Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.